# A short character sum in $\mathbb{F}_{p^{3}}$

### v1

@@ Abstract.
Homogeneous forms, Weil’s bound, short character sums, Multiplicative energy
@@ 1. Introduction
Short-character sums are of importance in analytic number theory. By "Short character sum", we mean that the summation is taken over an interval $I$ of size smaller than $q^{\frac{1}{2}}$ where $q$ is the modulus of the character. In our article characters are always multiplicative. Standard applications of these sums involve determining the smallest quadratic non-residue $(\bmod{q})$ and the distribution of primitive elements in finite fields. We shall revisit some history in this connection. Burgess’ bound [1] on Character Sums $\sum_{x\in I}\chi(x)$ of Dirichlet character in prime fields $\mathbb{F}_{p}$ provide a nontrivial estimate for an interval $I\subset[1,p-1]$ of size $|I|>p^{\frac{1}{4}+\varepsilon}$ for any $\varepsilon>0.$ Burgess result, which improves the Polya-Vinogradov inequality if $|I|<p^{1/2}$, was a major breakthrough. However, assuming the Riemann-Hypothesis we get a nontrivial upper bound as soon as $|I|>p^{\varepsilon}$, so there is a large space for improvement.
@@
provided that $|I_{i}|=p^{\rho_{d}+\varepsilon}$ for all $i=1,\ldots,d$ where $\rho_{d}=\frac{1}{2}-\frac{1}{2(d+1)}$. In addition, if one assumes that $\omega_{1},\ldots,w_{d}$ is a certain special type of basis, stronger results have been obtained by Burgess [1] and Karacuba [10]. Without such special assumptions, Mei-Chu Chang[11] improved on Davenport-Lewis(for $n\geq 5$) by obtaining a nontrivial upper bound for the intervals are of size $p^{\frac{2}{5}+\varepsilon}$. Furthermore, in dimension d=2, Chang [6] proved a nontrivial upper bound if $|I_{i}|=q^{1/4+\varepsilon}$ for any $\varepsilon>0$. In connection to this, we have our main theorem in $\mathbb{F}_{p^{3}}$.
@@
Given $\varepsilon>0$, there is a $\delta(\varepsilon)>0$ such that if $\omega\in\mathbb{F}_{p^{3}}-\mathbb{F}_{p}$ and $I,J$ are intervals of size $p^{3/8+\varepsilon}$($p$ sufficiently large), then
$\left|\sum_{x\in I\\ y\in J}\chi(x+\omega y)\right|<p^{-\delta(\varepsilon)}|I||J|$
for $\chi$ a nontrivial multiplicative character of $\mathbb{F}_{p^{3}}$.
@@
More generally, M.Chang[Theorem 10, [6]] established that if $\omega\in\mathbb{F}_{p^{d}}$ is a generator of $\mathbb{F}_{p}^{*}$ and if $I_{0},I_{1},\ldots,I_{d-1}$ are intervals of size $p^{\rho^{\prime}_{d}+\varepsilon}$, where $\rho^{\prime}_{d}=\frac{\sqrt{d^{2}+2d-7}+3-d}{8}$ then character sum over $\mathbb{F}_{p^{d}}$ is bounded by
@@
where the above $\delta(\varepsilon)>0$ is independent of the choice of $\omega$. Now, it is easy to check $\rho^{\prime}_{d}<\rho_{d}$. So, it is an improvement over the size of the intervals in Devanport-Lewis [7]. Our second main theorem is an extension of the previous Theorem to a subset of the basis. Unlike M.C. Chang’s theorem over the complete basis of $\mathbb{F}_{p^{d}}$, we consider character sums evaluated over sublattices of co-dimension 1 in $\mathbb{F}_{p^{d}}$. We prove the following
@@
$\sum_{x_{0}\in I_{0},\ldots,x_{d-2}\in I_{d-2}}\chi(x_{0}\omega^{n_{0}}+x_{1}\omega^{n_{1}}+x_{2}\omega^{n_{2}}+\ldots+x_{d-2}\omega^{n_{d-2}})<p^{-\delta(\varepsilon)}|J_{0}||J_{1}|\ldots|J_{d-2}|$
where, $n_{0}=0<n_{1}=1<\ldots<n_{d-2}<d$ and $\delta=\delta(\varepsilon)>0$ is independent of $\omega$.
Remarks: The novelty of our results is that we consider character sums for sublattices of codimension 1. This has not been studied in earlier work.
@@
Given $\varepsilon>0$, there is a $\delta(\varepsilon)>0$ such that the following holds: Let $\chi$ be a non-principle character in $\mathbb{F}_{p}$. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ which is not a perfect cube in moduli p. Let $I,J\subseteq[1,p-1]$ be interval of size
$|I|,|J|>p^{\frac{3}{8}+\varepsilon}$
@@
Let us recall some previous studies related to such sums. For $f(x,y)$ quadratic homogeneous polynomials, Burgess [4] proved the same type of estimate, where the interval length was $p^{\frac{1}{3}+\varepsilon}$ for large $p$. Chang [6] showed that if $f(x,y)$ is a quadratic form which is not a perfect square $\bmod{p}$, then there is a nontrivial upper bound if both intervals are of size $p^{\frac{1}{4}+\varepsilon}$. The above is application of M.C. Chang’s result [Theorem 11,[6]] to cubic binary forms. Previous to this work Burgess considered the multidimensional case [3],[4] in which
@@
is a product of n-linear forms over $\mathbb{Z}$ that are linearly independent over $\mathbb{F}_{p}$ where $p$ is a prime number and he proved that one can obtain a nontrivial upper bound of the character sum of the form if interval lengths are of same size $H=q^{\rho_{n}+\varepsilon}$. Bourgain and Chang incorporated ideas from Minkowski’s Subspace Theorem and geometry of numbers to get a nontrivial upper bound of the above character sum in a shorter range of $H$ which is $H=q^{\frac{1}{4}+\varepsilon}$, thus obtaining an n-dimensional result as strong as Burgess’s original threshold in each dimension[see,[5]]. More recently, L.Pierce and J.Xu [12] have established a Burgess type estimate for character sum evaluated on homogeneous forms of arbitrary degrees. In particular, they obtained a nontrivial upper bound for the character sum on the binary admissible forms if lengths of the intervals are of size $p^{\frac{1}{3}+\varepsilon}.$ However, our Theorem[3] provides a nontrivial upper bound of the character sum on the cubic binary if intervals are of size at least $q^{\frac{3}{8}+\varepsilon}$. Although the sizes of the intervals are slightly larger here compared to Pierce’s multidimensional result[12], we derive it by an alternative technique of "Multiplicative Energy of Set Tuples" in the context of character sum and by the use of Theorem[1]. In contrast to Pierce’s and Xu’s Stratification method[12] we shall use simple sum-product techniques to derive a Burgess type estimate for cubic binary forms at a cost of minor increment in the interval lengths. We shall introduce some definitions below to get a flavor of the additive combinatorial techniques used in our study.
Denote:
$E(A,I)=\{(x_{1},x_{2},t_{1},t_{2})\in A^{2}\times I^{2}:x_{1}t_{1}\equiv x_{2}t_{2}\}$
where $A\in\mathbb{F}_{p^{d}}$ is an arbitrary set and $I\subset[1,p-1]$ an interval. The underlying principle is actually related to Plunnecke-Ruza sum-set Theorey[8](the multiplicative version).
Acknowledgments. First and foremost, the author would like to express sincere gratitude to Prof. Stephan Baier for bringing this problem to his attention. The author thanks the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The research of the author was supported by a CSIR Ph.D. fellowship under file number 09/0934(13170)/2022-EMR-I.
@@ 2. Notations
Throughout this article, we will consider multiplying in the ring $R=\mathbb{F}_{p^{d}}$ or $\mathbb{F}_{p}\times\mathbb{F}_{p}$. We shall also denote $R^{*}$ by the group of invertible elements of the ring $R$. Let $\chi$ be a multiplicative character of $R$. Our notation is the following:
@@
    Note that Energy is always multiplicative in this article.
3. Multiplicative energy of two intervals in $\mathbb{F}_{p^{3}}$
Most of this calculations are analogous to M.C. Chang’s Paper[[6]] where it was explicitly calculated multiplicative energy of two intervals was explicitly calculated in $\mathbb{F}_{p^{2}}$. Here we shall calculate the energy of two intervals in $\mathbb{F}_{p^{3}}$.
@@ 3. Multiplicative energy of two intervals in $\mathbb{F}_{p^{3}}$
Let $\omega\in\mathbb{F}_{p^{3}}-\mathbb{F}_{p}$ and $Q=\{x+\omega y:x,y\in[1,\frac{1}{10}p^{\frac{3}{8}}]\}.$ Then for every $\eta\in\mathbb{F}_{p^{3}}$ the following holds:
|$\left|\left\{(z_{1},z_{2})\in Q\times Q:\eta=z_{1}.z_{2}\right\}\right]<exp\left(c\frac{\log p}{\log\log p}\right)$
@@
For given $\eta\in\mathbb{F}_{p^{3}}$, assume that $\eta$ can be written as a product of two elements in $Q$ in at least two ways. Let $x_{i}+\omega y_{i}\in Q$ such that
$(x_{1}+\omega y_{1})(x_{2}+\omega y_{2})=\eta=(x_{1}^{\prime}+\omega y_{1}^{\prime})(x_{2}^{\prime}+\omega y_{2}^{\prime}).$
So it implies $\omega$ satisfies a polynomial
$f(X)=(y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime})X^{2}+(x_{1}y_{2}+x_{2}y_{1}-x_{1}^{\prime}y_{2}^{\prime}-x_{2}^{\prime}y_{1}^{\prime})X+(x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime})$
in $\mathbb{Z}[X]$ in moduli $p$. But the minimal polynomial of $\omega$ is of degree $3$ so all the coefficient of the above polynomial is zero in the base field $\mathbb{F}_{p}.$ So, this implies that:
  1. (1)
    $(y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime})\equiv 0\bmod{p},$
  2. (2)
    $(x_{1}y_{2}+x_{2}y_{1}-x_{1}^{\prime}y_{2}^{\prime}-x_{2}^{\prime}y_{1}^{\prime})\equiv 0\bmod{p},$
  3. (3)
    $(x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime})\equiv 0\bmod{p}.$
From $(1),(2)$ and $(3)$ above and modulus value of these coefficients are bounded by $\frac{1}{25}p^{3/4}$ ,it follows the above inequality is true not only in $\mathbb{F}_{p}$ but also in $\mathbb{Z}$. Now from the divisorial bound of $p$, it is easy to see that
|$\left|\left\{(z_{1},z_{2})\in Q\times Q:\eta=z_{1}.z_{2}\right\}\right]\leq exp\left(c\frac{\log p}{\log\log p}\right).$
As an immediate consequence of Lemma $1$ are the following lemmas.
Lemma 5.
Let $Q$ be as in the above lemma, then the multiplicative energy $E(Q,Q)$ satisfies
$E(Q,Q)<exp\left(c\frac{\log p}{\log\log p}\right)|Q|^{2}.$
and
Lemma 6.
Let $Q$ be the same as before. Let $z_{1},z_{2}\in\mathbb{F}_{p^{3}}$, Then
$E(Q+z_{1},Q+z_{2})<exp\left(c\frac{\log p}{\log\log p}\right)|Q|^{2}.$
Proof.
See [Lemma 2 and 2’,[6]] ∎
4. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
Next we have a generalization of Lemma4.
Lemma 7.
Let $\omega\in\mathbb{F}_{p^{d}}$ be a generator over $\mathbb{F}_{p}$. Given $160<\sigma<11/$ 16 and let
$Q=$$\{x_{0}\omega^{n_{0}}+x_{1}\omega^{n_{1}}+x_{2}\omega^{n_{2}}+\ldots+x_{d-2}\omega^{n_{d-2}}:x_{i}\in[1,p^{\sigma}]$
$\forall i\in\{0,1,\ldots,d-2\}\text{ and }0\leq n_{0}<n_{1}<\ldots<n_{d-2}<d\}$
$Q_{0}=$$\{y_{0}+y_{1}\omega:y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\sigma}]\}$
Then
$\max_{\eta\in\mathbb{F}_{p}}|\{(z_{1},z_{2})\in Q\times Q_{0}:\eta=z_{1}z_{2}\}|<exp\left(c\frac{\log p}{\log\log p}\right).$
Proof.
The proof is similar to the proof of Lemma 4. ∎
Now we shall mention and prove a generalization of Lemma 5 and Lemma 6.
Lemma 8.
Let $Q,Q_{0}$ be as in Lemma 7 and let $z\in\mathbb{F}_{p^{d}}.$ Then
$E(z+Q,Q_{0})<exp\left(c_{d}\frac{\log p}{\log\log p}\right).|Q||Q_{1}|+|Q_{1}|^{2}.$
Proof.
The proof is similar to Lemma 5 and Lemma 6. ∎
@@ 4. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
Let $Q,Q_{0}$ be as in the previous lemmas with $|Q_{0}|<|Q|$ and let $I_{s}=[1,p^{1/k_{s}}]$ for $s=1,2,\ldots,r$ with $k_{s}\in\mathbb{Z}_{+}$ and $\sum\frac{1}{k_{s}}<1$. Then
$E(I_{1},I_{2},\ldots,I_{r},z+Q,Q_{0})<exp(c\frac{\log p}{\log\log p})|Q||Q_{0}|\prod_{s}|I_{s}|^{2(1-\sigma)}.$
@@
The proof is almost similar to Lemma 9 in [6], but we still need to mention the modifications that it required. First we start with the multiplicative energy:
$E(I_{1},I_{2},\ldots,I_{r},z+Q,Q_{0})=\frac{1}{p^{d}}\sum_{\chi}\prod\left|\sum_{t\in I_{s}}\chi(t)\right|^{2}\left|\sum_{\eta\in Q}\chi(z+\eta)\right|^{2}\left|\sum_{\eta\in Q_{0}}\chi(\eta)\right|^{2}.$
@@
(1) $\prod_{s=1}^{r}\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\}^{\frac{1}{k_{s}}}\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\}^{1-\sum\frac{1}{k_{s}}}.$
Notation:$\sum_{t\in I_{s}}\chi:=\sum_{t\in I_{s}}\chi(t),\sum_{\eta\in Q}\chi:=\sum_{\eta\in Q}\chi(z+\eta).$ For simplicity We denote
$A_{s}:=\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\}$
@@
$B:=\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\}.$
Now by Lemma 7 we get,
(2) $B=E(z+Q,Q_{0})<exp\left(c\frac{\log p}{\log\log p}\right)|Q||Q_{0}|.$

### v3

@@ Abstract.
Short Character Sum, Multiplicative Energy, Weil’s Bound, Binary Cubic Forms
@@ 1. Introduction
Short-character sums are of importance in analytic number theory. By "Short character sum", we mean that the summation is taken over an interval $I$ of size smaller than $q^{\frac{1}{2}}$ where $q$ is the modulus of the character. In our article, characters are always multiplicative, and intervals are intervals of integers, more precisely $a,b]:=\{n\in\mathbb{Z}:a\leq n\leq b\}$. Standard applications of these sums involve determining the smallest quadratic non-residue $(\bmod{q})$ and the distribution of primitive elements in finite fields. We shall revisit some history in this connection. Burgess’ bound [3] on Character Sums $\sum_{x\in I}\chi(x)$ of Dirichlet character in prime fields $\mathbb{F}_{p}$ provide a nontrivial estimate for an interval $I\subset[1,p-1]$ of size $|I|>p^{\frac{1}{4}+\varepsilon}$ for any $\varepsilon>0.$ Burgess result, which improves the Polya-Vinogradov inequality if $|I|<p^{1/2}$, was a major breakthrough. Assuming the Generalized Riemann Hypothesis (G.R.H), however, one can already obtain a non-trivial upper bound as soon as $|I|>p^{\varepsilon}$, leaving significant space for improvement in the unconditional case. Later, Granville and Soundararajan [13] showed that to achieve a non-trivial bound, one must have $|I|\gg(\log q)^{A}$ for any fixed positive $A$. In a subsequent paper [14], they further established that a non-trivial upper bound exists even under a weaker hypothesis—namely, that “100%” of the zeros of $L(s,\chi)$ up to height $1/4$ lie on the critical line.
@@
provided that $|I_{i}|=p^{\rho_{d}+\varepsilon}$ for all $i=1,\ldots,d$ where $\rho_{d}=\frac{1}{2}-\frac{1}{2(d+1)}$. In addition, if one assumes that $\omega_{1},\ldots,w_{d}$ is a certain special type of basis, stronger results have been obtained by Burgess [3] and Karacuba [17]. Without such special assumptions, Mei-Chu Chang [7] improved on the result of Davenport and Lewis (for $n\geq 5$) by obtaining a non-trivial upper bound for the size intervals $p^{\frac{2}{5}+\varepsilon}$. Furthermore, in dimension $d=2$, Chang [8] proved a non-trivial upper bound if $|I_{i}|=q^{1/4+\varepsilon}$ for any $\varepsilon>0$. Finally a generalization of Chang’s result to $\mathbb{F}_{p^{d}}$ and an improvement of Davenport and Lewis were established by Konyagin [18] for $|I_{i}|=q^{1/4+\varepsilon}$. In connection to these, we have our main theorem in $\mathbb{F}_{p^{3}}$.
@@
Let $\varepsilon>0$. Then there exists a constant $\delta=\delta(\varepsilon)>0$ such that for all sufficiently large primes $p$, the following holds. If $\omega\in\mathbb{F}_{p^{3}}\setminus\mathbb{F}_{p}$ and $I,J\subset\mathbb{F}_{p}$ are intervals of length at least $p^{3/8+\varepsilon}$, then for every nontrivial multiplicative character $\chi$ of $\mathbb{F}_{p^{3}}$,
$\left|\sum_{x\in I}\sum_{y\in J}\chi(x+\omega y)\right|\;\;<\;\;p^{-\delta(\varepsilon)}\,|I|\,|J|.$
@@
Remark 1: If we put $d=3$ and $\omega_{i}=\omega^{i-1}$ in the result of Davenport and Lewis then the length of the intervals becomes $\rho_{d}=\rho_{3}=3/8$. Unlike their theorem, our result allows one interval to be of smaller size.
Remark 2: For $\chi|{\mathbb{F}_{p}}\neq 1$ the theorem 1 follows from [12, Theorem]. However, our theorem also covers the case $\chi|_{\mathbb{F}_{p}}=1$ which was not encompassed in the previous results.
More generally, M.Chang [8, Theorem 10] established that if $\omega\in\mathbb{F}_{p^{d}}$ is a generator of $\mathbb{F}_{p}^{*}$ and if $I_{0},I_{1},\ldots,I_{d-1}$ are intervals of size $p^{\rho^{\prime}_{d}+\varepsilon}$, where $\rho^{\prime}_{d}=\frac{\sqrt{d^{2}+2d-7}+3-d}{8}$ then character sum over $\mathbb{F}_{p^{d}}$ is bounded by
@@
where the above $\delta(\varepsilon)>0$ is independent of the choice of $\omega$. Now, it is easy to check $\rho^{\prime}_{d}<\rho_{d}$. So, it is an improvement over the size of the intervals mentioned by Devanport and Lewis [9]. Our second main theorem is an extension of the previous Theorem to a subset of the basis. Unlike M. C. Chang’s theorem over the complete basis of $\mathbb{F}_{p^{d}}$, we consider character sums evaluated over sublattices of co-dimension 1 in $\mathbb{F}_{p^{d}}$. We prove the following
@@
$\sum_{x_{0}\in I_{0},\ldots,x_{d-2}\in I_{d-2}}\chi(x_{0}+x_{1}\omega+x_{2}\omega^{2}+\ldots+x_{d-2}\omega^{d-2})<p^{-\delta(\varepsilon)}|J_{0}||J_{1}|\ldots|J_{d-2}|$
where, $\delta=\delta(\varepsilon)>0$ is independent of $\omega$.
The novelty of our results is that we consider character sums for sublattices of codimension 1. This has not been studied in earlier work.
@@
Given $\varepsilon>0$, there is a $\delta(\varepsilon)>0$ such that the following holds: Let $\chi$ be a non-principle character in $\mathbb{F}_{p}$. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ which is non-degenerate over $\mathbb{F}_{p}$. Let $I,J\subseteq[1,p-1]$ be interval of size
$|I|,|J|=p^{\frac{3}{8}+\varepsilon}$
@@
Let us recall some previous studies related to such sums. For $f(x,y)$ quadratic homogeneous polynomials, Burgess [6] proved the same type of estimate, where the interval length was $p^{\frac{1}{3}+\varepsilon}$ for large $p$. Chang [8] showed that if $f(x,y)$ is a quadratic form which is not a perfect square $\bmod{p}$, then there is a nontrivial upper bound if both intervals are of size $p^{\frac{1}{4}+\varepsilon}$. The above is a generalization of M.C. Chang’s result [8, Theorem 11] to cubic binary forms. Recently, Heath-Brown[15] established that Chang’s result continues to hold for primitive characters even if $q$ is any odd square-free modulli with $(q,det(f))=1$. In a previous joint work with Baier [1, Proposition 5], the author proved that the above result also holds for non-principal characters. The present result is consistent with that earlier work.
Earlier, Burgess considered the multidimensional setup for this problem in [5], [6], where
@@
is a product of $n$ linear forms over $\mathbb{Z}$, which are linearly independent over $\mathbb{F}_{p}$, with $p$ a prime. He proved that a nontrivial upper bound for the corresponding character sum can be obtained when the interval lengths are all of the same size $H=q^{\rho_{n}+\varepsilon}$. Bourgain and Chang incorporated ideas from Minkowski’s Subspace Theorem and geometry of numbers to get a nontrivial upper bound of the above character sum in a shorter range of $H$ which is $H=q^{\frac{1}{4}+\varepsilon}$, thus obtaining a $n$-dimensional result as strong as Burgess’s original threshold in each dimension (see [2]0. More recently, L.Pierce and J.Xu [19] have established a Burgess type estimate for character sum evaluated on homogeneous forms of arbitrary degrees. In particular, they obtained a nontrivial upper bound for the character sum on the binary admissible forms if lengths of the intervals are of size $p^{\frac{1}{3}+\varepsilon}.$ However, our Theorem 3 provides a nontrivial upper bound of the character sum on the cubic binary if intervals are of size at least $q^{\frac{3}{8}+\varepsilon}$. Although the sizes of the intervals are slightly larger here compared to Pierce’s multidimensional result [19], we derive it by an alternative technique of "Multiplicative Energy of Set Tuples" in the context of character sum and by the use of Theorem [1]. In contrast to Pierce’s and Xu’s Stratification method [19] we shall use simple sum-product techniques to derive a Burgess type estimate for cubic binary forms at a cost of minor increment in the interval lengths. We shall introduce some definitions below to get a flavor of the additive combinatorial techniques used in our study.
Acknowledgments. The author expresses his sincere gratitude to Professor Stephan Baier for bringing this problem to his attention and for his valuable guidance. The author thanks the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The research of the author was supported by a CSIR Ph.D. fellowship under file number 09/0934(13170)/2022-EMR-I.
@@ 2. Notations
Throughout this article, we will consider multiplying in the ring $R=\mathbb{F}_{p^{d}}$ (multiplication defined by the multiplication of the field) or $\mathbb{F}_{p}\times\mathbb{F}_{p}$ (multiplication is defined coordinate-wise, $(a,b)\times(c,d)=(ac,bd)$. We shall also denote the group of invertible elements of the ring $R$ by $R^{*}$. Let $\chi$ be a multiplicative character of $R$. Our notation is the following:
@@
    Note that energy is always multiplicative in this article. Throughout the sequel, $c$ will be an arbitrarily small positive number which, depending on the context, may change from a line to the next. For $n\in\mathbb{N}$, we denote by $d(n)$ the number of divisors of $n$ and recall the well-known bound $d(n)\leq\exp\left(\frac{\log{n}}{\log\log{n}}\right)$.
3. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
First we shall obtain the following lemma which is required for calculating the multiplicative energy of a set triple.
@@ 3. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
Let $\omega\in\mathbb{F}_{p^{d}}$ be a generator over $\mathbb{F}_{p}$. Given $0<\sigma<\tfrac{11}{16}$, define
$Q=\left\{\,x_{0}+x_{1}\omega+x_{2}\omega^{2}+\cdots+x_{d-2}\omega^{d-2}\;:\;x_{i}\in[1,p^{\sigma}]\text{ for all }i\in\{0,1,\ldots,d-2\}\,\right\},$
and
$Q_{0}=\left\{\,y_{0}+y_{1}\omega\;:\;y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\sigma}]\,\right\}.$
Then, for all $\eta\in\mathbb{F}_{p^{d}}$,
$\max_{\eta\in\mathbb{F}_{p^{d}}}\bigl|\{\,(z_{1},z_{2})\in Q\times Q_{0}:\eta=z_{1}z_{2}\,\}\bigr|<\exp\!\left(c_{d}\frac{\log p}{\log\log p}\right),$
where $c_{d}>0$ is a constant depending only on $d$.
@@
Fix $\eta\in\mathbb{F}_{p^{d}}$ and suppose that $\eta$ can be represented as a product of two elements, one from $Q$ and one from $Q_{0}$, in at least two distinct ways. That is, assume
$(y_{0}+\omega y_{1}),(y_{0}^{\prime}+\omega y_{1}^{\prime})\in Q_{0},\quad(x_{0}+x_{1}\omega+\cdots+x_{d-2}\omega^{d-2}),(x_{0}^{\prime}+x_{1}^{\prime}\omega+\cdots+x_{d-2}^{\prime}\omega^{d-2})\in Q,$
such that
(1) $(x_{0}+x_{1}\omega+\cdots+x_{d-2}\omega^{d-2})(y_{0}+\omega y_{1})=(x_{0}^{\prime}+x_{1}^{\prime}\omega+\cdots+x_{d-2}^{\prime}\omega^{d-2})(y_{0}^{\prime}+\omega y_{1}^{\prime})=\eta.$
Hence, $\omega$ is a zero of the polynomial
$P(X)=(x_{0}+x_{1}X+\cdots+x_{d-2}X^{d-2})(y_{0}+Xy_{1})-(x_{0}^{\prime}+x_{1}^{\prime}X+\cdots+x_{d-2}^{\prime}X^{d-2})(y_{0}^{\prime}+Xy_{1}^{\prime})$
in $\mathbb{F}_{p}[X]$. Since the minimal polynomial of $\omega$ over $\mathbb{F}_{p}$ has degree $d$, it follows that all coefficients of $P(X)$ must vanish in $\mathbb{F}_{p}$. Moreover, since each coefficient is of magnitude strictly less than $p$, all these equalities also hold over $\mathbb{Z}$.
[…]
(2) $(x_{0}+x_{1}\widehat{\omega}+\cdots+x_{d-2}\widehat{\omega}^{d-2})(y_{0}+\widehat{\omega}y_{1})=(x_{0}^{\prime}+x_{1}^{\prime}\widehat{\omega}+\cdots+x_{d-2}^{\prime}\widehat{\omega}^{d-2})(y_{0}^{\prime}+\widehat{\omega}y_{1}^{\prime}).$
[…]
$\bigl|\{(z_{1},z_{2})\in\widehat{Q}\times\widehat{Q}_{0}:\widehat{\eta}=z_{1}z_{2}\}\bigr|<\exp\!\left(c\frac{\log p}{\log\log p}\right),$
where
$\widehat{Q}=\{\,x_{0}+x_{1}\widehat{\omega}+\cdots+x_{d-2}\widehat{\omega}^{d-2}:x_{i}\in[1,p^{\sigma}]\,\},\quad\widehat{Q}_{0}=\{\,y_{0}+y_{1}\widehat{\omega}:y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\sigma}]\,\}.$
This follows from the divisor bound for algebraic integers in $\mathbb{Z}[\widehat{\omega}$]. Indeed, for $\widehat{\zeta}\in\mathbb{Z}[\widehat{\omega}$],
$d_{Q(\widehat{\omega})}(\widehat{\zeta})^{2}=d\bigl(N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})\bigr)^{2}\leq\exp\!\left(2\,\frac{\log N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})}{\log\log N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})}\right)\leq\exp\!\left(2d^{2}(1+o(1))\,\frac{\log d\,\log p}{\log\log p}\right).$
Here, the middle inequality follows from the classical divisor estimate $d(n)\leq\exp\!\left(\frac{\log n}{\log\log n}\right),$ and the final bound arises from the size of the coefficients defining the elements of $\widehat{Q}$ and $\widehat{Q}_{0}$. ∎
The consequence here will be estimation of multiplicative energy of two intervals in $\mathbb{F}_{p^{d}}$.
Corollary 5.
Let $Q,Q_{0}$ be as in Lemma 4. Then
$E(Q,\,Q_{0})\;\leq\;\exp\!\Bigl(c_{d}\frac{\log p}{\log\log p}\Bigr)\,|Q|\,|Q_{0}|\;,$
@@
Let $Q,Q_{0}$ be as in the previous lemmas and let $I_{s}=[1,p^{1/k_{s}}]$ for $s=1,2,\ldots,r$ with $k_{s}\in\mathbb{Z}_{+}$ and $\sum\frac{1}{k_{s}}<1$. Then
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})\ \leq\exp(c\frac{\log p}{\log\log p})|Q||Q_{0}|\prod_{s}|I_{s}|^{\frac{1}{10}(19-4\sigma)}.$
@@
The proof is similar to [8, Lemma 9], but we still give a proof here for the sake of completeness. First we start with the multiplicative energy of intervals:
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})=\frac{1}{p^{d}}\sum_{\chi}\prod\left|\sum_{t\in I_{s}}\chi(t)\right|^{2}\left|\sum_{\eta\in Q}\chi(\eta)\right|^{2}\left|\sum_{\eta\in Q_{0}}\chi(\eta)\right|^{2}.$
@@
(3) $\prod_{s=1}^{r}\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}^{\frac{1}{k_{s}}}\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}^{1-\sum\frac{1}{k_{s}}}.$
Notation: $\sum_{t\in I_{s}}\chi:=\sum_{t\in I_{s}}\chi(t),\sum_{\eta\in Q}\chi:=\sum_{\eta\in Q}\chi(\eta).$ For simplicity we denote
$A_{s}:=\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}$
@@
$B:=\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}.$
Now by Corrolary 5 we get,
(4) $B=E(Q,Q_{0})\leq\exp\left(c\frac{\log p}{\log\log p}\right)|Q||Q_{0}|.$